karl is buying a new car. At bank A, a car loan will cost him 11.5%. At bank B, it will cost 12.25%. If he borrows $8,000, how much will he save if he borrows from bank A?
Answer by Maths68(1474)

Difference in %age =12.25-11.5=0.75%
0.75% of $8,000 = (.75*8000)/100=60
He will save $60
Another way to solve this problem
12.25% of 8000 = (12.25*8000)/100=980
11.50% of 8000 = (11.5*8000)/100=920
At bank B he will have to pay extra = 980
At bank A he will have to pay extra = 920
if he borrows money from Bank A he will save = 980-920=$60